OpenStudy (anonymous):

Fiona is serving iced tea and lemonade at a picnic. She has only 44 cups in which to serve the drinks. If x represents the number of glasses of iced tea and y represents the number of glasses of lemonade, which equation represents the number of drinks she can serve? x – y = 44 44 + y = x y – x = 44 x = 44 – y

OpenStudy (alexalas):

remember x+y=44

OpenStudy (anonymous):

but thats not an answer choice ..

OpenStudy (alexalas):

okay say x=22 and y=22 which equation up top is equal?

OpenStudy (anonymous):

im still not sure

OpenStudy (alexalas):

remember x+y=44 x+y-y=44-y x=44-y

OpenStudy (anonymous):

oh okay makes sense now. thanks

OpenStudy (alexalas):

any more question?

OpenStudy (anonymous):

not as of now, no

OpenStudy (alexalas):

okay bye